OpenStudy (raffle_snaffle):

Let y = sales at a fast-food outlet (1000s of $), x1 = number of competing outlets within a 1-mile radius, x2 = population within a 1-mile radius (1000s of people), and x3 be an indicator variable that equals 1 if the outlet has a drive-up window and 0 otherwise. Suppose that the true regression model is Y = 8.00 − 1.5x1 + 6.9x2 + 15.1x3 + ε (a) What is the mean value of sales when the number of competing outlets is 3, there are 7000 people within a 1-mile radius, and the outlet has a drive-up window? (Round your answer to one decimal place.) (b) What is the mean value of sales for an

OpenStudy (raffle_snaffle):

(b) What is the mean value of sales for an outlet without a drive-up window that has two competing outlets and 5000 people within a 1-mile radius? (Round your answer to one decimal place.)

OpenStudy (raffle_snaffle):

@hartnn

OpenStudy (raffle_snaffle):

@hartnn

hartnn (hartnn):

you know how to find the mean value with Y that is given?

OpenStudy (raffle_snaffle):

i plugged those numbers in and got a number

OpenStudy (raffle_snaffle):

48318.60

hartnn (hartnn):

ε = ..?

OpenStudy (raffle_snaffle):

What is epsalon?

OpenStudy (raffle_snaffle):

I had a problem similar to this one and didn't plug a value in for epsolan

hartnn (hartnn):

ok, so just assumed ε = 0

OpenStudy (raffle_snaffle):

48318.60

OpenStudy (raffle_snaffle):

you get the same? It's wrong

hartnn (hartnn):

i got the same, but is Y the actual mean? can you post a screenshot on how the other similar problem was solved?

OpenStudy (raffle_snaffle):

yeah give me a sec.

OpenStudy (raffle_snaffle):

hold on it's taking forever to post my attachment

OpenStudy (raffle_snaffle):

let me log off and back on...

OpenStudy (raffle_snaffle):

I think the wifi i am using is bad. I can't post it.

hartnn (hartnn):

no problem, can you just describe how they did it?

OpenStudy (raffle_snaffle):

Cardiorespiratory fitness is widely recognized as a major component of overall physical well-being. Direct measurement of maximal oxygen uptake (VO2max) is the single best measure of such fitness, but direct measurement is time-consuming and expensive. It is therefore desirable to have a prediction equation for VO2max in terms of easily obtained quantities. Consider the variables y = VO2max (L/min) x1 = weight (kg) x2 = age (yr) x3 = time necessary to walk 1 mile (min) x4 = heart rate at the end of the walk (beats/min) Here is one possible model for male students, consistent with the information given in the article "Validation of the Rockport Fitness Walking Test in College Males and Females"†: Y = 5.0 + 0.01x1 − 0.05x2 − 0.13x3 − 0.01x4 + ε σ = 0.4

OpenStudy (raffle_snaffle):

1.73

OpenStudy (raffle_snaffle):

(b) What is the expected value of VO2max when weight is 79 kg, age is 18 yr, walk time is 14 min, and heart rate is 134 b/m?

OpenStudy (raffle_snaffle):

There is the question and my answer.

OpenStudy (raffle_snaffle):

@Michele_Laino